OpenStudy (anonymous):

I need to turn this (x-24)(x-18) into a quadratic equation using reverse foil method. PLEASE HELP

OpenStudy (anonymous):

It's not the reverse FOIL method... It's just FOIL. So you multiply First, Outside, Inside, and Last.

OpenStudy (anonymous):

thatt's what I thought but we have to work our problem like this: go from this: (x – 5)(x – 18) = 0 to this: 2x2 – 40x + 200 = 6x + 20

OpenStudy (lgbasallote):

reverse FOIL? so LIOF?

OpenStudy (anonymous):

I guess....i'm totally confused

hero (hero):

(x-24)(x-18) = x(x-18) -24(x-18) = x^2 - 18x - 24x + 432 = x^2 - 42x + 432

OpenStudy (anonymous):

Okay so first, is x*x outside is x*-18 inside is x*-5 last is -5*-18 Combine all that and you get this: \[x^2 - 18x - 5x + 90\] Simplified: \[x^2 - 23x + 90\]

OpenStudy (jiteshmeghwal9):

& put a zero @ the R.H.S LOL;)

OpenStudy (anonymous):

I would have, of course, but I think fee only needed the equation... not to solve it

OpenStudy (anonymous):

ok how do you go from (x-5)(x-18)=0 to 2x^2 – 40x + 200 = 6x + 20

OpenStudy (jiteshmeghwal9):

\[\LARGE{Hint:-(a+b)(c+d)=a(c+d)+b(c+d)}\]

OpenStudy (anonymous):

in this formula what is d?

OpenStudy (jiteshmeghwal9):

a,b,c,d are assumed as constants . here u only assume a=x b=24 c=x d=18

OpenStudy (jiteshmeghwal9):

& this equation also can simplified by the identity\[\LARGE{(x+a)(x+b)=x^2+(a+b)x+ab}\]

OpenStudy (jiteshmeghwal9):

here a=24 & b=18.

OpenStudy (jiteshmeghwal9):

x remains same according to ur equation.

OpenStudy (jiteshmeghwal9):

gt it ?? can u do it @Fee ?

OpenStudy (anonymous):

so after plugging everything in it should look like this: (x+24)(x+18)=x^2+(24+18)x+24*18

OpenStudy (jiteshmeghwal9):

yup :)

OpenStudy (jiteshmeghwal9):

x^2+42x+432.

OpenStudy (jiteshmeghwal9):

ohh ! u don't look the sign it's not positive here it's negative therefore it will look like x^2-(24+18)x+24*18 hence we gt x^2-42x+432.

OpenStudy (jiteshmeghwal9):

\[\LARGE{(x-a)(x-b)=x^2-(a+b)x+ab}\]

hero (hero):

What was wrong with my methods?

OpenStudy (anonymous):

@Hero nothing....same thing I get all those steps but in the example we were given it started off with 2 numbers: 5 and 18

OpenStudy (anonymous):

from there t shows (x-5)(x-18)

OpenStudy (jiteshmeghwal9):

actually i'm only teaching him other methods but in this question he is said to solve the question through FOIL method so ur method isn't wrong @Hero :)

OpenStudy (anonymous):

then from there it says apply rules of algebra to arrive at this: 2x^2 – 40x + 200 = 6x + 20 this is where I get lost

hero (hero):

(x – 5)(x – 18) = `x(x-18) -5(x - 18) = x^2 - 18x - 5x + 90` = x^2 - 23x + 90 = 0

OpenStudy (anonymous):

then we have to factor back down to the starting point

hero (hero):

2x^2 -40x + 200 = 0 x^2 - 20x + 100 = 0 x^2 - 10x - 10x + 100 = 0 x(x - 10) - 10(x-10) = 0 (x - 10)(x - 10) = 0 (x - 10)^2 = 0

hero (hero):

x - 10 = 0 x = 10

hero (hero):

If you're still confused, I can show you step by step on vyew